Circular No. 8999

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V679 CARINAE = NOVA CARINAE 2008 E. O. Waagen and A. Henden, AAVSO, report the discovery of an ap-

parent nova (mag 7.9) by Katarzyna Malek, Center for Theoretical Physics of the Polish Academy of Sciences, on an image taken on Nov. 26.26 UT with the “Pi of the Sky” automated survey (original report apparently via vsnet-alert 10740). The object, which was at mag 7.55 on Nov. 27.27, was reported by Malek to be located at α = 11^{h}13^{m}54^{s}, δ = −61^{o}14′00′′ (equinox 2000.0). A. Maury, S. Barnes, and C. Harlingten have measured the position of the nova on unfiltered CCD images taken on Dec. 1.266 with a 50-cm f/6.9 reflector located at San Pedro de Atacama, Chile: α = 11^{h}13^{m}53^{s}.78, δ = −61^{o}13′48′′.1, magnitude 7.4. N. Samus and E. V. Kazarovets report that the GCVS team assigns the designation V679 Car to this object.

H. E. Bond, Space Telescope Science Institute, writes that a spectrogram (range 365–545 nm; resolution 0.43 nm) of V679 Car was obtained with the SMARTS 1.5-m telescope on Cerro Tololo by M. Hernandez and F. Walter on Nov. 30.37 UT. The spectrum confirms that the object is a classical nova, of the “Fe II” type. The spectrum is dominated by broad emission lines of the Balmer series and of Fe II emission features. The FWZI of Hβ is ∼ 4000 km/s. Continued spectroscopic and photometric observations are urged.

Visual magnitude estimates for V679 Car: Nov. 29.120 UT, 8.3 (A. Amorim, Florianopolis, Brazil); 30.775, 8.4 (N. J. Brown, Ellenbrook, West Australia). CCD magnitude estimates for V679 Car by D. Hanzl, San Pedro de Atacama, Chile: Nov. 25.317, 8.6 (Canon 20Da camera, R channel); 28.313, 7.9; Dec. 2.241, 8.9 (Canon 300D camera, R). Nothing is visible at this location on Hanzl’s images from Nov. 22.258–22.266 (limiting mag 12.0) and 24.268–24.301 (limiting mag 13.9).

P. Schmeer, Bischmisheim, Germany, forwards the following position for V679 Car from C. Gualdoni: α = 11^{h}13^{m}53^{s}.79, δ = −61^{o}13′48′′.2 (equinox 2000.0); Schmeer adds that this is only 1′′.4 from a BMW-Chandra source located at position end figures 53^{s}.66, 49′′.3.

COMETS 206P/BARNARD-BOATTINI AND 207P/NEAT Comet P/2008 T3 (Barnard-Boattini) = P/1892 T1 (cf. IAUC 8995)

has been given the permanent designation 206P. Comet P/2008 T5 (NEAT) = P/2001 J1 (cf. IAUC 8996) has been given the permanent designation 207P. See also MPC 64001.

2008 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green